# StreamDiffusion: A Pipeline-level Solution for Real-time Interactive Generation

### v1

@@ Abstract
$*{}^{*}$ denotes equal contributionThis work was done when Toshiki was a remote intern at UC Berkeley
We introduce StreamDiffusion, a real-time diffusion pipeline designed for interactive image generation. Existing diffusion models are adept at creating images from text or image prompts, yet they often fall short in real-time interaction. This limitation becomes particularly evident in scenarios involving continuous input, such as Metaverse, video game graphics rendering, live video streaming, and broadcasting, where high throughput is imperative. To address this, we present a novel and simple approach that transforms the original sequential denoising into the batching denoising process. Stream Batch eliminates the conventional wait-and-interact approach and enables fluid and high throughput streams. To handle the disparity in frequencies between data input and model throughput, we design a novel input-output queue for parallelizing the streaming process. Moreover, the existing diffusion pipeline uses classifier-free guidance to enforce the generated results conditioned on the prompts. We point out that current implementations are inefficient due to the inherent redundancy of negative conditional denoising. To mitigate the redundant computations, we propose a novel residual classifier-free guidance (RCFG) algorithm that reduces the number of negative conditional denoising steps to only one or even zero. Besides, to optimize power consumption, we use a simple and effective stochastic similarity filtering strategy, which significantly reduces GPU activation frequency, thereby improving GPU utilization efficiency. Our batching denoising strategy achieves around 1.5x speedup compared to the sequential denoising method at different denoising levels. The proposed (RCFG) leads to up-to 2.05x speedups compared to the previous conventional classifier-free guidance. The combination of the proposed strategies and existing mature acceleration tools makes the image-to-image generation achieve up-to 91.07fps on one RTX 4090 GPU, improving the throughputs of AutoPipline developed by Diffusers over 59.56x. Beyond the pipeline efficiency, our proposed StreamDiffusion also significantly reduces the energy consumption by 2.39x on one RTX 3060 GPU and 1.99x on one RTX 4090 GPU for a static scene input, respectively. The code is available at https://github.com/cumulo-autumn/StreamDiffusion.
Figure 1: Real-time image-to-image results from camera and screen captures. The first and second columns show examples of AI-assisted drawing in real-time, and the third column displays the real-time rendering of 2D illustrations from 3D avatars. The fourth and fifth columns demonstrate the real-time camera filter.
@@ 1 Introduction
Recently, there has been a growing trend in the commercialization of diffusion models [21, 23, 3] for applications within the Metaverse, as well as in the realms of online video streaming and broadcasting. These areas require diffusion pipelines that offer high throughput and low latency to ensure efficient human interaction. A pertinent example is the use of diffusion models to create virtual YouTubers. These digital personas should be capable of reacting in a fluid and responsive manner to user input.
To advance high throughput and real-time interactive capabilities, current efforts primarily focus on reducing the number of denoising iterations, such as decreasing from 50 iterations to just a few [16, 17] or even one [28, 13]. The common strategy is to distill the multi-step diffusion models into a few steps or re-frame the diffusion process with neural Ordinary Differential Equations (ODE) [14, 26]. Quantization has also been applied to diffusion models [12, 9] to improve efficiency.
@@
In this paper, we start from an orthogonal direction and introduce StreamDiffusion, a pipeline-level solution that enables real-time interactive image generation with high throughput. We highlight that existing model design efforts can still integrate with our pipeline. Our approach enables the use of N-step denoising diffusion models while still keeping high throughput and offers users more flexibility in choosing their preferred models.
We leverage a simple strategy in StreamDiffusion: instead of the original sequential denoising mode, we batch the denoising steps. We are inspired by the asynchronous processing in the computer architecture pipeline, and point out that we do not need to wait for the previous denoising processes to finish then to start the current denoising step, as shown in Fig. 2. In order to deal with the synchronous issue of input frequency and U-Net processing frequency, we leverage a queue strategy to cache the inputs and output. Our pipeline is different from merely asynchronous processing. We make use of the advantage of GPU parallelism; thus, we can use one U-Net to denoise a batched noise latent feature. More importantly, the existing diffusion pipeline incorporated classifier-free guidance to emphasize the given prompts in the generated images. Nevertheless, traditional implementations of classifier-free guidance are plagued by excessive and redundant computational overheads. We introduce an innovative approach termed as residual classifier-free guidance (RCFG). The novel technique approximates the negative condition with a virtual residual noise, which allows us to calculate the negative condition noise only during the initial step of the process.
@@
Further, we point out that maintaining the diffusion models always in an interactive mode is energy-consuming as it keeps hitting GPU. To reduce the energy, we further apply a stochastic similarity filtering strategy. In the pipeline, we compute the similarities between continuous inputs and determine whether the diffusion model should process the images based on the similarity.
@@
Experiments demonstrate that our proposed StreamDiffusion can achieve up to 91.07fps for image generation on a one RTX4090 GPU, surpassing the diffusion Autopipeline from Diffusers team by up to 59.6x. Besides, our stochastic similarity filtering strategy significantly reduces the GPU power usage by 2.39x on one RTX 3090GPU and by 1.99x on one RTX 4090GPU. Our proposed StreamDiffusion is a new diffusion pipeline that is not only efficient but also energy-saving.
2 Related works
2.1 Diffusion Models
@@
Diffusion models, as introduced in [25, 8], have sparked considerable interest in the commercial sector due to their high-quality image generation capabilities and the level of control they offer over the generated images. These models have been progressively adapted for various applications, including text-to-image generation [20, 19, 2], image edition [1, 22], and video generation [4, 5].
A notable advancement in this area is the development of consistent models [26], which have demonstrated the potential to improve the efficiency of the sampling process with rarely compromising on image quality. Building on this, recent studies have focused on enhancing the efficiency of diffusion models by reducing the number of sampling steps required, further expanding their practical applicability and efficiency.
2.2 Accelerating Diffusion Model
Diffusion models are currently limited by their slow speed in generating outputs. To address this drawback, several strategies have been developed. These include methods that don’t require additional training, like using Ordinary Differential Equation (ODE) solvers, as explored by [26, 14, 15, 17]. Other techniques involve adaptive step size solvers as discussed by [10] and predictor-corrector methods noted by [25]. The training-based methods are also proposed, such as optimized discretization techniques [27], neural operator design [29], and model distillation [24]. Our proposed method distinguishes itself significantly from the approaches mentioned previously. While earlier methods primarily focus on the low latency of their individual model designs, our approach takes a different route. We introduce a comprehensive pipeline-level solution specifically tailored for high throughputs, which enables efficient interactive diffusion. Our pipeline is designed to seamlessly integrate most of the low-latency diffusion models discussed above. By doing so, we achieve a notable increase in speed within the interactive generation process.
3 StreamDiffusion Pipeline
StreamDiffusion is a new diffusion pipeline aiming for high throughput. It comprises several key components: the Stream Batch strategy, Residual Classifier-Free Guidance (RCFG), an input-output queue, Stochastic Similarity Filter, a pre-computation procedure, and model acceleration tools with a tiny-autoencoder. We elaborate on the details below.
Figure 3: Virtual residual noise vectors: The orange vectors depict the virtual residual noise that starts from the PF ODE trajectory and points to the original input latent $x_{0}$
3.1 Batching the denoise step
In diffusion models, denoising steps are performed sequentially, resulting in a proportional increase in the processing time of U-Net relative to the number of steps. However, to generate high-fidelity images, it is necessary to increase the number of steps. To resolve the high-latency generation in interactive diffusion, we propose a method called Stream Batch.
@@ 3.2 Residual Classifier-Free Guidance
Classifier-free guidance (CFG) [7] is an algorithm that enhances the effect of the original conditioning by performing vector calculations between an unconditioning or a negative conditioning term [6] and the original conditioning term (Eq. 1).
@@
where $\epsilon_{\tau_{i},c}$ and $\epsilon_{\tau_{i},\bar{c}}$ are predicted residual noise at the denoising step $\tau_{i}$ with conditioning embedding $c$ and negative conditioning embedding $\bar{c}$, respectively, and $\gamma$ is a guidance scale.
This allows for benefits such as strengthening the effect of the prompt. However, to compute the negative conditioning residual noise, it is necessary to pair each input latent variable with the negative conditioning embedding and pass it through the U-Net at every inference time. To address this drawback, we introduce Residual Classifier-Free Guidance (RCFG), which significantly reduces the computational cost of the additional U-Net inference for the negative conditioning embedding.
Firstly, the encoded input latent $x_{0}$ can be transferred to the noise distribution $x_{\tau_{0}}$ as follows,
where $\alpha_{\tau_{0}}$ and $\beta_{\tau_{0}}$ are values determined by a noise scheduler and $\epsilon_{0}$ is a sampled noise from a Gaussian $\mathcal{N}(0,I)$. Following the implementation of latent consistency model [26], we can predict the data distribution $\hat{x}_{0,{\tau_{i-1}},\mathrm{cfg}}$ and generate a next step noise distribution $x_{\tau_{i},\mathrm{cfg}}$ using the CFG residual noise $\epsilon_{{\tau_{i-1}},\mathrm{cfg}}$ as follows,
$\hat{x}_{0,{\tau_{i-1}},\mathrm{cfg}}=\frac{x_{{\tau_{i-1}},\mathrm{cfg}}-\sqrt{\beta_{\tau_{i-1}}}\epsilon_{{\tau_{i-1}},\mathrm{cfg}}}{\sqrt{\alpha_{\tau_{i-1}}}}$ (3)
$x_{{\tau_{i}},\mathrm{cfg}}=\sqrt{\alpha_{\tau_{i}}}\hat{x}_{0,{\tau_{i-1}},\mathrm{cfg}}+\sqrt{\beta_{\tau_{i}}}\epsilon_{i},$ (4)
Where ${\epsilon_{i}}\sim\mathcal{N}(0,I)$.
For the next step CFG calculation, instead of computing the next step negative conditioning residual noise $\epsilon_{{\tau_{i}},\bar{c}}$, which requires an additional U-Net computation cost, we hypothesize the existence of virtual residual noise $\epsilon_{{\tau_{i}},\bar{c}^{\prime}}$, and a virtual negative condition embedding $\bar{c}^{\prime}$ that predicts the original input image’s latent $x_{0}$ from $x_{{\tau_{i}},\mathrm{cfg}}$, as shown in Fig. 3. Similarly to Eq. 3, this virtual residual noise satisfies Eq. 5 to predict the input latent $x_{0}$.
$x_{0}\approx\hat{x}_{0,{\tau_{i}},\bar{c}^{\prime}}=\frac{x_{{\tau_{i}},\mathrm{cfg}}-\sqrt{\beta_{\tau_{i}}}\epsilon_{{\tau_{i}},\bar{c}^{\prime}}}{\sqrt{\alpha_{\tau_{i}}}}$ (5)
Given the initial value $x_{0}$, and the subsequent values of $x_{{\tau_{i}},\mathrm{cfg}}$ obtained sequentially through Eq. 4, the virtual residual noise $\epsilon_{{\tau_{i}},\bar{c}^{\prime}}$ can be analytically determined by employing these values with the Eq. 5:
$\epsilon_{{\tau_{i}},\bar{c}^{\prime}}=\frac{x_{{\tau_{i}},\mathrm{cfg}}-\sqrt{\alpha_{\tau_{i}}}x_{0}}{\sqrt{\beta_{\tau_{i}}}}$ (6)
With the virtual residual noise $\epsilon_{{\tau_{i}},\bar{c}^{\prime}}$ obtained from Eq. 6, we formulate RCFG as follows,
@@
RCFG that uses the original input image latent $x_{0}$ as the residual term can effectively generate results that diverge from the original input image according to the magnitude of the guidance scale $\gamma$, thereby enhancing the effect of conditioning without the need for additional U-Net computations. We call this method Self-Negative RCFG.
Not only to deviate from the original input image $x_{0}$, but also to diverge from any negative condition, we can compute the negative conditioning residual noise $\epsilon_{\tau_{0},\bar{c}}$ using U-Net only onetime for the first denoising step and use this to predict the negative conditioning denoised data as follows,
@@
In Eq. 6, instead of $x_{0}$, using $\hat{x}_{0,{\tau_{0}},\bar{c}}$, we can obtain the virtual residual noise $\epsilon_{{\tau_{i+1}},\bar{c}^{\prime}}$ that can effectively diverge the generation results from the controllable negative conditioning embedding $\bar{c}$. We call this Onetime-Negative RCFG.
By combining Eq. 2, Eq. 6, and Eq. 8, both Self-Negative RCFG and Onetime-Negative RCFG can be uniformly represented by the following equation,
$\epsilon_{{\tau_{i}},\bar{c}^{\prime}}=\frac{x_{{\tau_{i}},\mathrm{cfg}}-\sqrt{\alpha_{\tau_{i}}}x_{0}}{\sqrt{\beta_{\tau_{i}}}}+\sqrt{\frac{\alpha_{\tau_{i}}\beta_{\tau_{0}}}{\beta_{\tau_{i}}\alpha_{\tau_{0}}}}(\epsilon_{\tau_{0},\bar{c}}-\epsilon_{0})$ (9)
When $\epsilon_{\tau_{0},\bar{c}}=\epsilon_{0}$, Eq. 9 is identical to Eq. 6 and returns Self-Negative RCFG result.
In contrast to the conventional CFG, which requires $2n$ computations of U-Net, the Self-Negative RCFG and Onetime-Negative RCFG necessitate only $n$ and $n+1$ computations of U-Net, respectively, where $n$ is the number of the denoising steps.
Figure 5: Inference pipeline overview: The core diffusion inference pipeline, including VAE and U-Net. By incorporating a denoising batch and pre-computed prompt embedding cache, sampled noise cache, and scheduler values cache, the inference pipeline is enhanced in terms of speed and enables real-time image generation. Stochastic Similarity Filter (SSF) is designed to save GPU power usage. It dynamically gates the pass of the diffusion model. This framework realizes fast and energy-efficient real-time inference.
3.3 Input-Output Queue

### v2

@@ Abstract
denotes equal contributionThis work was done when Toshiki was a remote intern at UC Berkeley
We introduce StreamDiffusion, a real-time diffusion pipeline designed for streaming image generation. Existing diffusion models are adept at creating images from text or image prompts, yet they often fall short in real-time interaction. This limitation becomes particularly evident in scenarios involving continuous input, such as augmented/virtual reality, video game graphics rendering, live video streaming, and broadcasting, where high throughput is imperative. StreamDiffusion tackles this challenge through a novel pipeline-level system design. It employs unique strategies like batching the denoising process (Stream Batch), residual classifier-free guidance (R-CFG), and stochastic similarity filtering (SSF). Additionally, it seamlessly integrates advanced acceleration technologies for maximum efficiency. Specifically, Stream Batch reformulates the denoising process by eliminating the traditional wait-and-execute approach and utilizing a batching denoising approach, facilitating fluid and high-throughput streams. This results in 1.5x higher throughput compared to the conventional sequential denoising approach. R-CFG significantly addresses inefficiencies caused by repetitive computations during denoising. It optimizes the process to require minimal or no additional computations, leading to speed improvements of up to 2.05x compared to previous classifier-free methods. Besides, our stochastic similarity filtering dramatically lowers GPU activation frequency by halting computations for static image flows, achieving a remarkable reduction in computational consumption—2.39 times on an RTX 3060 GPU and 1.99 times on an RTX 4090 GPU, respectively. The synergy of our proposed strategies with established acceleration technologies enables image generation to reach speeds of up to 91.07 fps on a single RTX 4090 GPU, outperforming the throughput of AutoPipeline, developed by Diffusers, by more than 59.56x.
Figure 1: StreamDiT-4B: video generation can be streaming and real-time.
@@ 1 Introduction
Recently, there has been a growing trend in the commercialization of diffusion models [32, 34, 3, 28] for applications within the entertainment industry such as Metaverse, online video streaming, broadcasting, and even the robotic field [7]. A pertinent example is the use of diffusion models to create virtual YouTubers. These digital personas should be capable of reacting in a fluid and responsive manner to user input. These areas require diffusion pipelines that offer high throughput and low latency to ensure the efficient interactive streaming generation.
@@
To advance the efficiency, current efforts primarily focus on reducing the number of denoising steps, such as decreasing from 50 denoise steps to just a few [24, 25] or even one [42, 21]. The strategy includes distilling the multi-step diffusion models into a few steps [36, 40] or re-framing the diffusion process with neural Ordinary Differential Equations (ODE) [22, 23]. Quantization has also been applied to diffusion models [18, 14] to improve efficiency. These methods share the common goal of approximating either the diffusion process itself or the original model weights to achieve efficiency gains. In this paper, we aim at an orthogonal direction and introduce StreamDiffusion, a pipeline-level solution that enables streaming image generation with high throughput. We highlight that existing model design efforts can still be integrated with our pipeline. Our approach enables the use of N-step denoising diffusion models while keeping high throughput and offers users more flexibility in choosing their preferred models.
Figure 2: The overview of StreamDiffusion. StreamDiffusion combines several key components: (1) Stream Batch efficiently processes the denoising steps in batches. (2) Residual Classifier-Free Guidance approximates the negative condition term to reduces the unnecessary repetitive calculations in the UNet. (3) Stochastic Similarity Filter controls the pass of the image stream by calculating the similarity between frames to eliminate the redundant hit onto GPUs. Furthermore, StreamDiffusion leverages techniques like input-output queues for smooth data flow, cache management for faster processing through pre-calculated embedding, and a tiny VAE model to further contribute to overall efficiency. This synergistic combination allows StreamDiffusion to generate high-quality images at high throughputs while consuming minimal energy.
Specifically, StreamDiffusion seamlessly incorporates a suite of novel strategies. Among these, we propose a simple yet novel approach termed Stream Batch. This method differs from the traditional sequential denoising mode, instead of batching the denoising steps. This subtle modification enhances efficiency without sacrificing the quality of image generation. We highlight Stream Batch enables a new capability of generating images conditioned on future frames in a streaming mode, which is something impossible in previous works. Via injecting the future frames, Stream Batch significantly improves the temporal consistency with few additional overhead. Furthermore, the key novelty of Stream Batch lies not only in its GPU parallelization; rather, it serves as a practical realization of a broader Stream Denoising framework. By denoising inputs at diagonally-offset timesteps within a streaming queue structure, diffusion models naturally achieve continuous, autoregressive generation—emitting one output frame per each newly sampled frame—while benefiting from parallel computation. Crucially, this design generalizes directly to sequential tasks such as video, audio, or robotic action-sequence generation, enabling interactive, unbounded-length synthesis.
Besides, we point out that it is time-consuming for existing diffusion pipelines to use classifier-free guidance for emphasizing the prompts during generation, due to the repetitive and redundant computations for negative conditions. To address this issue, we introduce an innovative approach termed as residual classifier-free guidance (R-CFG). R-CFG approximates the negative condition with a virtual residual noise, which allows us to calculate the negative condition noise only during the initial step of the process. We also indicate that using the original input image latent as the residual term effectively generates results that diverge from the original input image according to the magnitude of the guidance scale, which is a special case of our R-CFG and does not require any computations for the negative condition term.
@@
Furthermore, in real applications such as virtual youtuber and AR/VR cases, maintaining the diffusion models always in an active mode is energy-consuming as it keeps hitting GPU. To reduce the energy, we further apply a stochastic similarity filtering (SSF) strategy. In the pipeline, we compute the similarities between continuous inputs and determine whether the diffusion model should process the images based on the probability of an activated similarity. This enables both energy efficiency and visual fluency. In order to further improve the efficiency to cater to the real applications, we apply simple yet effective engineering implementations such as Input-Output Queue (IO-Queue), pre-computing caching, and TensorRT. The overview of the StreamDiffusion pipeline is shown in Fig. 2.
@@
Experiments demonstrate that our proposed StreamDiffusion can achieve up to 91.07fps for image generation on one RTX4090 GPU, surpassing the diffusion Autopipeline from Diffusers [15] team by up to 59.6x. Besides, our stochastic similarity filtering strategy significantly reduces the GPU power usage by 2.39x on one RTX 3090GPU and by 1.99x on one RTX 4090GPU. Our proposed StreamDiffusion is a new diffusion pipeline that is not only efficient but also energy-saving.
2 Related work
Efficient Diffusion Models
Diffusion models [39, 13, 32, 29] have sparked considerable interest in the commercial sector due to their high-quality image/video generation capabilities. These models have been progressively adapted for various applications, including text-to-image generation [31, 30, 2], image editing [1, 33], video generation [4, 5] and even perception [41, 19, 16]. However, diffusion models are currently limited by their slow speed in generating outputs. In response to this challenge, a variety of strategies have been proposed. One of the mainstreams is to approximate the SDE-based diffusion process [39, 38] through an ordinary differentiable equation (ODE) framework. For example, DPM and DPM++ [22, 23] introduce ODE-based samplers, which significantly reduce the hundreds of denoising steps to between 15 and 20. Building upon the ODE formulation, InstaFlow [21] advances the reduction of denoising steps to a single instance through the novel strategy of rectified flow [20], while achieving performance close to that of Stable Diffusion [32]. Additionally, distillation from pre-trained diffusion models has also been explored as a method to facilitate few-step denoising. For instance, the consistency model [40] leverages the principle of self-consistency between noise at different denoising steps and uses pre-trained diffusion models [32] to guide the learning of a few-step denoising model, thereby enabling the generation of images within a minimal number of steps. In a notable extension of this concept, LCM [25, 26] applies the idea to the latent space rather than the pixel space. The use of distillation methods [36, 35, 42] to enhance the efficiency of the original Stable Diffusion model presents promising results. Besides improving efficiency through the lens of reducing denoising steps, quantization methods [18, 14] are proposed to make the model run in the regime of low float points, albeit with the potential trade-off of fidelity and efficiency. Moreover, parallel sampling [37] tries to utilize the approximate-parallel denoising strategy to improve the latency of the diffusion model. We emphasize that our work focuses on enhancing throughput and our work is orthogonal to [37]. Notably, our method is optimized for single-GPU, which are common among users. In contrast, the parallel sampling method reduces throughput on a single GPU.
Our proposed StreamDiffusion is significantly different from the approaches mentioned previously. While earlier methods primarily focus on the low latency of their individual model designs, our approach takes a different route. We introduce a pipeline-level solution specifically tailored for high throughput. Our pipeline can seamlessly integrate the low-latency diffusion models discussed above. Our proposed Stream Batch, residual classifier-free guidance, and the integration of other efficiency-enhancement methods focus on improving the efficiency of the whole pipeline instead of a single diffusion model.
@@ 2 Related work
Classifier-free Guidance
Classifier-free guidance [12] is widely used for conditional generation due to the simplicity, efficiency, and stability compared to classifier-guidance [9]. It leverages negative prompts [8, 11, 32] and essentially operates the vector arithmetic shift in latent space, i.e., we take a step of size (usually set by the guidance scale) away from the unconditional vector or negatively conditioned vector in the direction toward the conditioning manifold [12]. In the practical implementation, the classifier-free guidance is conducted by sharing the same UNet for both the conditional term and unconditional (or negative) term and subtracting the effect of the unconditional term from the conditioned one. We point out that the way of multiple denoising processes leads to unnecessary computations for the unconditional (or negative) term. To get rid of these redundant computations, we propose a novel residual-classifier-free guidance, termed as R-CFG, which approximates the conditional noise prediction with only requiring one or even zero-time computation for the negatively conditioned noise prediction for the UNet. We note that R-CFG is especially designed for the SDEdit method [27], as we mainly focus on the applications of translating streaming image flow. Our proposed R-CFG significantly improves the latency of the conditional image-to-image generation.
3 StreamDiffusion
StreamDiffusion is a new diffusion pipeline aiming for high throughput. It comprises three key components: Stream Batch strategy, Residual Classifier-Free Guidance (R-CFG), and Stochastic Similarity Filter. Besides, we also incorporate other acceleration methods like a novel input-output queue designed by us, the pre-computation procedure, the tiny-autoencoder, and model acceleration tools such as TensorRT. We elaborate on the details below.
Figure 4: Virtual residual noise vectors: The orange vectors depict the virtual residual noise that starts from PF ODE trajectory and points to the original input latent $x_{0}$
3.1 Stream Batch: Batching the Denoise Step
In diffusion models, denoising steps are performed sequentially, resulting in a proportional increase in the processing time of U-Net relative to the number of steps. However, to generate high-fidelity images, it is necessary to increase the number of steps. To resolve this problem in interactive diffusion, we propose a method called Stream Batch.
@@ 3.1 Stream Batch: Batching the Denoise Step
Waiting and Batching can also increase the throughput of the diffusion pipeline. However, with naive Waiting and Batching (WB), denoising cannot begin immediately on the first input frame, leading to higher latency compared to Stream Batch. We mainly aim for smooth streaming applications. Yet achieving a smooth frame rate with WB requires additional engineering, such as precise inference speed estimation and input-output frame synchronization, and minor timing errors must be carefully managed. In contrast, Stream Batch automatically ensures a consistent interval between input and output frames, providing the advantage of lower latency while dynamically reaching the optimal throughput.
3.2 Improve Time Consistency by Stream Batch
Maintaining temporal consistency in video generation is challenging. Many approaches ensure frame coherence by referencing past frames, often through cross-frame attention. However, our Stream Batch method uniquely enables temporal consistency using information from future frames. As shown in Fig. 3, Stream Batch allows simultaneous denoising of multiple frames, passing information from future frames to the current frame. This supports real-time image translation that adapts to sudden changes in input while preserving consistency. In Stream Batch with $n$ denoising steps, keys and values for each frame at each time step form the following batches:
$K_{\text{Batch}}=\left[K_{t+i,0},\dots,K_{t,i},\dots,K_{t-(n-1-i),n-1}\right]$
$V_{\text{Batch}}=\left[V_{t+i,0},\dots,V_{t,i},\dots,V_{t-(n-1-i),n-1}\right]$
These key and value batches incorporate information across different time frames and denoising steps. For example, if the frame at time step $t$ has reached the $i$-th denoising step, the batch includes $i$ future denoising steps for different frames and $n-1-i$ past frames. In Stream Batch Cross-frame Attention, rather than using the typical $K_{t,i}$ and $V_{t,i}$, we employ $K_{\text{Batch}}$ and $V_{\text{Batch}}$, which integrate past and future frame information for the attention computation:
$\text{Attn}(Q_{t,i},K_{\text{Batch}},V_{\text{Batch}})=\text{Softmax}\left(\frac{Q_{t,i}\cdot K_{\text{Batch}}^{T}}{\sqrt{d}}\right)V_{\text{Batch}}$
This approach enables the generation process to account for information from both past and future frames, as well as across different denoising stages, thus enhancing temporal consistency.
@@ 3.3 Residual Classifier-Free Guidance
Firstly, SDEdit based method [27] adds perturbation to the input image $x_{0}$ and transfers it to the noise distribution $x_{\tau_{0}}$ as follows,
where $\alpha_{\tau_{0}}$ and $\beta_{\tau_{0}}$ are values determined by a noise scheduler and $\epsilon_{0}$ is a sampled noise from a Gaussian $\mathcal{N}(0,I)$. When using consistency models for conditional image editing, $x_{\tau_{0}}$ can be considered as a point on the PF ODE trajectory, which leads to the conditioning manifold. To intensify the conditioning by Classifier-Free Guidance (CFG)[10], it is imperative to compute a noise for a negative condition PF ODE trajectory, which is used in vector arithmetic shift for the guidance (Eq. 2).
@@
This requirement introduces additional computational overhead at each denoising step. To reduce this computational overhead, R-CFG utilizes the fact that the original input image $x_{0}$ is referable at any stage of denoising steps.
For any latent $x_{\tau_{i}}$ at the denoising step $\tau_{i}$, we can assume the existence of the virtual negative condition $\bar{c}_{\tau_{i}}^{\prime}$, that satisfies the self-consistency described as Eq. 3 This implies that $x_{\tau_{i}}$ is on the PF ODE trajectory going back to the input image $x_{0}$.
$x_{0}\approx\hat{x}_{0,{\tau_{i}},\bar{c}_{\tau_{i}}^{\prime}}=f_{\theta}(x_{{\tau_{i}}},\tau_{i},\bar{c}_{\tau_{i}}^{\prime})$ (3)
Following the LCM model parameterization [25] and our approximation for the inference time skip connections ($c_{\mathrm{skip}}(\tau)=0$, $c_{\mathrm{out}}(\tau)=1$ at $\tau\neq 0$), the self-consistency equation (Eq. 3) can be expressed as follows,
$x_{0}\approx\frac{x_{{\tau_{i}}}-\sqrt{\beta_{\tau_{i}}}\epsilon_{{\tau_{i}},\bar{c}_{\tau_{i}}^{\prime}}}{\sqrt{\alpha_{\tau_{i}}}}$ (4)
Given the initial value $x_{0}$, and the subsequent values of $x_{{\tau_{i}}}$ obtained sequentially through the iterative denoising, the virtual noise vector $\epsilon_{{\tau_{i}},\bar{c}^{\prime}}$ in the direction toward the input image can be analytically determined by employing these values with the Eq. 4:
$\epsilon_{{\tau_{i}},\bar{c}^{\prime}}=\frac{x_{{\tau_{i}}}-\sqrt{\alpha_{\tau_{i}}}x_{0}}{\sqrt{\beta_{\tau_{i}}}}$ (5)
With the virtual noise $\epsilon_{{\tau_{i}},\bar{c}^{\prime}}$ obtained from Eq. 5, we formulate R-CFG by:
@@
R-CFG that uses the original input image latent $x_{0}$ as the residual term can effectively generate results that diverge from the original input image according to the magnitude of the guidance scale $\gamma$, thereby enhancing the effect of conditioning without the need for additional U-Net computations. We call this method Self-Negative R-CFG.
Not only to deviate from the original input image $x_{0}$, but also to diverge from any negative condition, we can find the desired reference point $\hat{x}_{0,{\tau_{0}},\bar{c}}$ that reflects the negative condition $\bar{c}$ using the same self-consistency formulation:
@@
We can obtain $\hat{x}_{0,{\tau_{0}},\bar{c}}$ by computing the actual negative conditioned noise $\epsilon_{\tau_{0},\bar{c}}$ using U-Net only one time for the first denoising step.
In Eq. 5, instead of $x_{0}$, using $\hat{x}_{0,{\tau_{0}},\bar{c}}$, we can obtain the virtual negative conditioned noise $\epsilon_{{\tau_{i+1}},\bar{c}^{\prime}}$ that can effectively diverge the generation results from the controllable negative conditioning $\bar{c}$. We name this method Onetime-Negative R-CFG. In contrast to the conventional CFG, which requires $2n$ computations of U-Net, the Self-Negative RCFG and Onetime-Negative RCFG necessitate only $n$ and $n+1$ computations of U-Net, respectively, where $n$ is the number of the denoising steps.